# Mitigating Metropolitan Carbon Emissions with Dynamic Eco-driving at Scale

### v1

@@ 2 Introduction
For interrupted traffic flow dynamics, dynamic eco-driving has attracted significant attention in recent decades Mintsis et al. (2020). Dynamic eco-driving is concerned with reducing carbon emissions and/or fuel consumption of vehicles individually and at the fleet level by strategically modifying driving patterns and behaviors. (Figure 1). Owing to advances in vehicle-to-infrastructure (V2I) communication technologies and the standardization of the signal phase and timing (SPaT) messages, these technologies have facilitated the research of dynamic eco-driving, especially near signalized intersections, where traffic flow is persistently interrupted Mintsis et al. (2020). Based on prior studies, we crudely estimate that persistent yet unproductive idling and excess accelerations at signalized intersections could contribute up to 14.6% of land transportation CO2 emissions in the United States (US) (refer to Appendix 10.1). Highlighting its significance, this is comparable to half of the well-scrutinized US airline sector emissions Lee et al. (2009). Reducing emissions at signalized intersections thus emerges as a promising avenue for achieving climate change mitigation goals.
@@
However, despite nearly four decades of research, large-scale implementation of dynamic eco-driving is yet to be seen. A potential reason and a notable observation is that existing studies vary widely in assumptions and, thus, reported outcomes. In particular, previous studies report a significant variation in emission benefits ranging from 2% to 56% Mintsis et al. (2020). Furthermore, our observation is that a majority of studies are small-scale (e.g., one or two intersections or a few intersections in a route) and mainly focus on technical questions that concern how to eco-drive Yang et al. (2016, 2020); Sun et al. (2020). But the fundamental socio-technical question of whether to eco-drive, which requires going beyond a few intersections-based analyses, remains unanswered Barkenbus (2010). Our aim of this work is to contribute a large-scale prospective analysis of eco-driving at signalized intersections, highlighting city-scale emission benefits and uncovering insights that may influence widespread adoption—factors that are often elusive in smaller-scale studies. Such an analysis provides a scientific basis for informed decision-making, strategic planning of infrastructure, and technology development Barkenbus (2010).
However, conducting a large-scale analysis of eco-driving presents significant challenges, a complexity that has largely eluded in previous studies. Diverse eco-driving scenarios stemming from factors such as intersection topologies, vehicle-related factors (e.g., fuel type, age, model, etc.), weather conditions, and seasonal changes must be considered to ensure that the prospective eco-driving impact assessments are representative Falkenauer (1998); Jayawardana et al. (2022). It not only ensures that insights are generalizable but also aligns with industry best practices for impact assessments iee (2022); Ross and Guhathakurta (2017); Duinker and Greig (2007). While field studies are considered a gold standard, as they reflect real-world conditions, it is prohibitively expensive to conduct traffic field tests at scale Gloudemans et al. (2023). Simulation-based analyses offer greater scalability, and uncertainties stemming from the sim-to-real gap can be addressed Vinitsky et al. (2022); Jang et al. (2019).
@@
We identified 33 influential factors for eco-driving based on previous studies and expert opinion. An eco-driving scenario manifests as a combination of these factor valuations. Each factor can take multiple valuations, creating a vast scenario space of at least $2^{33}$ scenarios. A comprehensive eco-driving analysis thus suffers from the curse of dimensionality Donoho et al. (2000). The high-dimensional scenario space grows exponentially with each additional factor, amplifying computational complexity in scenario modeling and eco-driving control. Furthermore, in eco-driving control, each scenario involves solving an optimal eco-driving control problem, making a comprehensive large-scale analysis challenging due to the need to solve a vast number of them—a scale far out of reach of previous studies.
@@
This article addresses these challenges to comprehensively analyze eco-driving at signalized intersections at scale. We consider semi-autonomous vehicles, such as vehicles equipped with advanced driver assistance systems (ADAS), which we can be leveraged to intelligently manage stop-and-go traffic, effectively reducing carbon emissions. To navigate the complexity of this undertaking, we take a data-driven approach to scenario modeling and model the scenario space with representative scenarios instead of exhausting all scenarios with inputs from domain experts. We then optimize eco-driving control across this––still significantly large––spectrum of representative scenarios by employing model-free multi-task deep reinforcement learning and zero-shot transfer learning instead of often used model-based methods. To make the optimization of eco-driving control tractable, we employ a carefully designed network decomposition strategy. This approach focuses on optimizing eco-driving at each intersection individually, instead of network scope optimizations, which is a significantly hard problem and intractable even with advances in modern computing and thus remains an open optimization challenge Qadri et al. (2020). While network decomposition strategies are commonly used in eco-driving research, many studies fail to analyze the impact of eco-driving optimized on individual intersections when deployed at the network scale. For example, if the throughput is increased due to eco-driving at an optimized intersection, that may lead to a queue spillback event in the following intersection at the network level. This oversight can render the previous analyses less reliable, potentially leading to network-level traffic overflows when such strategies are implemented. In contrast, we ensure that such impacts are accounted for in the assessment to ensure that the benefits observed in our analysis are consistent and reliable at the network level.
@@
Analyzing nearly one million traffic scenarios stemming at 6,110 intersections across three major high-emissions US metropolitan cities Gately et al. (2015)- Atlanta, Los Angeles, and San Francisco- our analysis reveals insights that were elusive at small scale analyses and strengthen the existing knowledge with high confidence. We find that implementing eco-driving could reduce US intersection CO2 emissions by 11-22%, which under reasonable extrapolating assumptions is equivalent to the national carbon emissions of Israel and Nigeria (Refer to Appendix 10.2). Such an implementation does not affect intersection throughput and could even assist in increasing throughput due to smoothed traffic flow. Furthermore, we observe that eco-driving is as safe as conventional driving, measured by safety surrogate measures Wang et al. (2021a). With just 10% of eco-driving adoption, 25%-50% of the total emission benefits (benefits if the adoption is 100%) can be achieved, underscoring its significant near-term potential as a climate change mitigation strategy. However, when eco-driving adoption increases, the impact level of each factor affecting eco-driving benefits changes. For instance, while longer green signal times may negatively affect benefits at low adoption levels, they can positively contribute at higher adoption rates. This dynamic relationship necessitates frequent adjustments in traffic engineering practices and a reevaluation of long-term infrastructure investments at intersections. Furthermore, approximately 70% of the total benefits are concentrated in just 20% of intersections, even at lower adoption levels. However, the composition of this high-impact subset of intersections varies considerably across different adoption levels, with minimal overlap. This finding underscores the need for a more nuanced and adaptive approach to eco-driving deployment, emphasizing the importance of careful planning and strategic implementation to maximize benefits as adoption rates evolve over time.
@@ 3.1 Eco-driving at signalized intersections
Over the past decade, numerous studies have investigated the challenge of dynamic eco-driving at signalized intersections. While they often study different traffic settings, Mintsis et al. Mintsis et al. (2020) show an average eco-driving benefit of 17%, with a standard deviation of 11% by considering 21 past studies. While this highlights the potential of eco-driving, these studies often consider only a few factors that are known to have an impact on eco-driving benefits, making the shown benefits scenario/traffic setting specific. Based on the literature, we identify nearly 33 factors that affect eco-driving benefits. We refer the reader to Appendix 10.3 for more details on these factors and their known impact.
Moreover, most eco-driving studies focus on devising better eco-driving strategies (e.g., how to eco-drive). Such studies often take three forms: 1) analytical approaches, 2) simulation approaches, and 3) field experiments Mintsis et al. (2020). Analytical approaches often focus on analyzing the eco-driving problem from a theoretical point of view by analyzing properties such as the stability of the system under certain eco-driving strategies Coppola et al. (2022). On the other hand, simulation-based approaches are used to test the potential of new eco-driving strategies and also to conduct impact assessments. However, one of the major limitations of such work is the lack of consideration of many factors that affect eco-driving benefitsBarkenbus (2010); Mintsis et al. (2020); Jayawardana and Wu (2022); Meng and Cassandras (2020). Furthermore, for simplicity and tractability of eco-driving control optimization scope, most of these studies analyze a few intersections under simplified traffic conditions (e.g., uniform inflow rates) Yang et al. (2016, 2020); Sun et al. (2020). A few studies that focus on multiple intersections often solely focus on the generalization of eco-driving strategies and rely on synthetically generated traffic scenarios while not considering the impact assessment requirements Jayawardana et al. (2024). Finally, field studies are also conducted to demonstrate the benefits of eco-driving and to validate the applicability of eco-driving strategies Hao et al. (2018); Stahlmann et al. (2018); Mintsis et al. (2017). However, these studies are often confined to small regions (e.g., a few intersections) due to the logistical difficulties of conducting such experiments at scale Mintsis et al. (2020).
@@ 3.2 Traffic intersection scenario modeling
Numerous studies have focused on modeling intersections for various purposes, such as traffic signal control Ault and Sharon (2021) and autonomous driving Leurent (2018). However, they often remain limited to specific scenarios Wang et al. (2021b); Ault and Sharon (2021), primarily due to a focus on proof of concepts rather than comprehensive assessments of proposed methods. Additionally, modeling multiple intersections can be prohibitively expensive, requiring data fusion from various sources and handling missing data.
However, our requirements for modeling intersection scenarios are objectively and strategically different from previous work, necessitating us to develop our own datasets. First, we are neither presenting an eco-driving proof of concept nor attempting to answer how to do eco-driving. Our goal is to analyze the impact of eco-driving to answer the question of whether to eco-drive. Such questions can not be answered by analyzing merely a few intersection scenarios. Unless we have a representative set of intersection scenarios, merely relying on a select intersection could lead to evaluation overfitting where the insights extracted from a few scenarios do not generalize to the class of scenarios Jayawardana et al. (2022); Whiteson et al. (2011).
Furthermore, we must avoid overfitting to a specific region in our impact assessment Gately et al. (2015). While there are popular intersection datasets available for Manhattan Island in New York Wei et al. (2019), using them could lead to regional overfitting. Manhattan Island exhibits distinct spatial and temporal traffic dynamics due to its block-oriented road network. Insights drawn from an impact assessment solely based on these intersections would not be effective in generalizing to other cities or regions.
Another option is to use synthetic datasets, which are easier to generate Li et al. (2022). However, impact assessments based on these datasets are not particularly meaningful as they may not reflect the real-world intersection distributions and may contain oversimplified scenarios Feng et al. (2023).
Thus, the fundamental requirement for modeling intersection scenarios lies in their diversity, encompassing variations at both the intersection and regional levels and making sure they are drawn from the real-world distribution. However, it is crucial to exercise caution when creating datasets, ensuring that the fused data originate from the same distribution. For instance, the IntD dataset Bock et al. (2020) containing a set of intersections in Germany is popular within the research community, but combining it with U.S.-specific data like vehicle inflows may result in a distribution mismatch.
@@ 3.3 Methods for eco-driving control
Earlier studies on eco-driving have predominantly explored five control strategies for controlling vehicles. They include 1) dynamic programming, 2) heuristics, 3) trajectory optimizations, 4) model-based reinforcement learning (model-based RL), and 5) model-free reinforcement learning (model-free RL). Each of these approaches exhibits distinct advantages and drawbacks. In Table 1, we offer a comparative analysis of these methods, focusing on five prevalent shortcomings.
@@
Dynamic programming[1,2] x x x x
Heuristics[3] x x
Trajectory optimization[4,5] x x * *
Model-based RL[6,7] * * x *
Model-free RL[8-10] * *
Table 1: A comparison of the shortcomings of five known methodologies for eco-driving control policy synthesis. The ’✓’ indicates that the method does not demonstrate the shortcoming, an ’x’ indicates that the method demonstrates the shortcoming, and a ’*’ indicates that the shortcoming may or may not present. [1]Maamria et al. (2016), [2]Doan et al. (2018), [3]Katsaros et al. (2011), [4]Kamal et al. (2010), [5]Wang et al. (2019), [6]Lee et al. (2020), [7]Zhu et al. (2022), [8]Jayawardana and Wu (2022), [9]Wegener et al. (2021), [10]Bai et al. (2022)
  • Sub optimality: Refers to the method’s ability to produce optimal results.
@@
  • High computational load: Indicates the run-time computational burden needed for obtaining solutions. This is crucial, considering that vehicles may not have powerful onboard computers.
  • Small scale: Refers to the optimization problem’s scale, e.g., one intersection or two. It does not address deployment scale directly, although some synergies may exist.
@@ 3.4 Reinforcement learning for vehicle-based traffic control
Contrary to conventional traffic control mechanisms such as traffic signals and stop signs, vehicle-based traffic control is increasingly being explored in recent research. Such vehicle-based control is referred to as Lagrangian traffic control and involves controlling a fleet of vehicles to achieve certain fleet-level (including both controlled and uncontrolled vehicles) objectives such as congestion management Vinitsky et al. (2018), reducing emissions Jayawardana and Wu (2022), and fostering smoother traffic flow Wu et al. (2021); Yan et al. (2022). While various studies have presented proofs-of-concept, they still fall short of accurately modeling real-world scenarios Yan et al. (2022); Wu et al. (2017); Di and Shi (2021) or go beyond toy traffic problems Yan et al. (2022).
Existing research often relies on simplified traffic models, such as toy traffic blocks Yan et al. (2022), to assess the benefits of Lagrangian traffic control methods. However, none of these studies explore the potential of large-scale implementations that could offer meaningful insights for practical decision-making. While toy tasks are valuable for gaining preliminary insights into a method, they might not resonate strongly with policymakers and stakeholders due to the lack of rigorous analysis of the benefits at a scale that holds significance Esser and Rickert (2000).

### v2

@@ 2 Introduction
For interrupted traffic flow dynamics, dynamic eco-driving has attracted significant attention in recent decades (38). Dynamic eco-driving is concerned with reducing carbon emissions and/or fuel consumption of vehicles individually and at the fleet level by strategically modifying driving patterns and behaviors. (Figure 1). Owing to advances in vehicle-to-infrastructure (V2I) communication technologies and the standardization of the signal phase and timing (SPaT) messages, these technologies have facilitated the research of dynamic eco-driving, especially near signalized intersections, where traffic flow is persistently interrupted (38). Based on prior studies, we crudely estimate that persistent yet unproductive idling and excess accelerations at signalized intersections could contribute up to 14.6% of land transportation CO2 emissions in the United States (US) (refer to Appendix 10.1). Highlighting its significance, this is comparable to half of the well-scrutinized US airline sector emissions (29). Reducing emissions at signalized intersections thus emerges as a promising avenue for achieving climate change mitigation goals.
@@
However, despite nearly four decades of research, large-scale implementation of dynamic eco-driving is yet to be seen. A potential reason and a notable observation is that existing studies vary widely in assumptions and, thus, reported outcomes. In particular, previous studies report a significant variation in emission benefits ranging from 2% to 56% (38). Furthermore, our observation is that a majority of studies are small-scale (e.g., one or two intersections or a few intersections in a route) and mainly focus on technical questions that concern how to eco-drive (64; 63; 46). But the fundamental socio-technical question of whether to eco-drive, which requires going beyond a few intersections-based analyses, remains unanswered (4). Our aim of this work is to contribute a large-scale prospective analysis of eco-driving at signalized intersections, highlighting city-scale emission benefits and uncovering insights that may influence widespread adoption—factors that are often elusive in smaller-scale studies. Such an analysis provides a scientific basis for informed decision-making, strategic planning of infrastructure, and technology development (4).
However, conducting a large-scale analysis of eco-driving presents significant challenges, a complexity that has largely eluded in previous studies. Diverse eco-driving scenarios stemming from factors such as intersection topologies, vehicle-related factors (e.g., fuel type, age, model, etc.), weather conditions, and seasonal changes must be considered to ensure that the prospective eco-driving impact assessments are representative (15; 24). It not only ensures that insights are generalizable but also aligns with industry best practices for impact assessments (20; 41; 12). While field studies are considered a gold standard, as they reflect real-world conditions, it is prohibitively expensive to conduct traffic field tests at scale (18). Simulation-based analyses offer greater scalability, and uncertainties stemming from the sim-to-real gap can be addressed (50; 21).
@@
We identified 33 influential factors for eco-driving based on previous studies and expert opinion. An eco-driving scenario manifests as a combination of these factor valuations. Each factor can take multiple valuations, creating a vast scenario space of at least $2^{33}$ scenarios. A comprehensive eco-driving analysis thus suffers from the curse of dimensionality (11). The high-dimensional scenario space grows exponentially with each additional factor, amplifying computational complexity in scenario modeling and eco-driving control. Furthermore, in eco-driving control, each scenario involves solving an optimal eco-driving control problem, making a comprehensive large-scale analysis challenging due to the need to solve a vast number of them—a scale far out of reach of previous studies.
@@
This article addresses these challenges to comprehensively analyze eco-driving at signalized intersections at scale. We consider semi-autonomous vehicles, such as vehicles equipped with advanced driver assistance systems (ADAS), which we can be leveraged to intelligently manage stop-and-go traffic, effectively reducing carbon emissions. To navigate the complexity of this undertaking, we take a data-driven approach to scenario modeling and model the scenario space with representative scenarios instead of exhausting all scenarios with inputs from domain experts. We then optimize eco-driving control across this––still significantly large––spectrum of representative scenarios by employing model-free multi-task deep reinforcement learning and zero-shot transfer learning instead of often used model-based methods. To make the optimization of eco-driving control tractable, we employ a carefully designed network decomposition strategy. This approach focuses on optimizing eco-driving at each intersection individually, instead of network scope optimizations, which is a significantly hard problem and intractable even with advances in modern computing and thus remains an open optimization challenge (39). While network decomposition strategies are commonly used in eco-driving research, many studies fail to analyze the impact of eco-driving optimized on individual intersections when deployed at the network scale. For example, if the throughput is increased due to eco-driving at an optimized intersection, that may lead to a queue spillback event in the following intersection at the network level. This oversight can render the previous analyses less reliable, potentially leading to network-level traffic overflows when such strategies are implemented. In contrast, we ensure that such impacts are accounted for in the assessment to ensure that the benefits observed in our analysis are consistent and reliable at the network level.
@@
Analyzing nearly one million traffic scenarios stemming at 6,110 intersections across three major high-emissions US metropolitan cities (17)- Atlanta, Los Angeles, and San Francisco- our analysis reveals insights that were elusive at small scale analyses and strengthen the existing knowledge with high confidence. We find that implementing eco-driving could reduce US intersection CO2 emissions by 11-22%, which under reasonable extrapolating assumptions is equivalent to the national carbon emissions of Israel and Nigeria (Refer to Appendix 10.2). Such an implementation does not affect intersection throughput and could even assist in increasing throughput due to smoothed traffic flow. Furthermore, we observe that eco-driving is as safe as conventional driving, measured by safety surrogate measures (52). With just 10% of eco-driving adoption, 25%-50% of the total emission benefits (benefits if the adoption is 100%) can be achieved, underscoring its significant near-term potential as a climate change mitigation strategy. However, when eco-driving adoption increases, the impact level of each factor affecting eco-driving benefits changes. For instance, while longer green signal times may negatively affect benefits at low adoption levels, they can positively contribute at higher adoption rates. This dynamic relationship necessitates frequent adjustments in traffic engineering practices and a reevaluation of long-term infrastructure investments at intersections. Furthermore, approximately 70% of the total benefits are concentrated in just 20% of intersections, even at lower adoption levels. However, the composition of this high-impact subset of intersections varies considerably across different adoption levels, with minimal overlap. This finding underscores the need for a more nuanced and adaptive approach to eco-driving deployment, emphasizing the importance of careful planning and strategic implementation to maximize benefits as adoption rates evolve over time.
@@ 3.1 Eco-driving at signalized intersections
Over the past decade, numerous studies have investigated the challenge of dynamic eco-driving at signalized intersections. While they often study different traffic settings, (38) show an average eco-driving benefit of 17%, with a standard deviation of 11% by considering 21 past studies. While this highlights the potential of eco-driving, these studies often consider only a few factors that are known to have an impact on eco-driving benefits, making the shown benefits scenario/traffic setting specific. Based on the literature, we identify nearly 33 factors that affect eco-driving benefits. We refer the reader to Appendix 10.3 for more details on these factors and their known impact.
Moreover, most eco-driving studies focus on devising better eco-driving strategies (e.g., how to eco-drive). Such studies often take three forms: 1) analytical approaches, 2) simulation approaches, and 3) field experiments (38). Analytical approaches often focus on analyzing the eco-driving problem from a theoretical point of view by analyzing properties such as the stability of the system under certain eco-driving strategies (8). On the other hand, simulation-based approaches are used to test the potential of new eco-driving strategies and also to conduct impact assessments. However, one of the major limitations of such work is the lack of consideration of many factors that affect eco-driving benefits(4; 38; 25; 36). Furthermore, for simplicity and tractability of eco-driving control optimization scope, most of these studies analyze a few intersections under simplified traffic conditions (e.g., uniform inflow rates) (64; 63; 46). A few studies that focus on multiple intersections often solely focus on the generalization of eco-driving strategies and rely on synthetically generated traffic scenarios while not considering the impact assessment requirements (23). Finally, field studies are also conducted to demonstrate the benefits of eco-driving and to validate the applicability of eco-driving strategies (19; 45; 37). However, these studies are often confined to small regions (e.g., a few intersections) due to the logistical difficulties of conducting such experiments at scale (38).
@@ 3.2 Traffic intersection scenario modeling
Numerous studies have focused on modeling intersections for various purposes, such as traffic signal control (2) and autonomous driving (31). However, they often remain limited to specific scenarios (54; 2), primarily due to a focus on proof of concepts rather than comprehensive assessments of proposed methods. Additionally, modeling multiple intersections can be prohibitively expensive, requiring data fusion from various sources and handling missing data.
However, our requirements for modeling intersection scenarios are objectively and strategically different from previous work, necessitating us to develop our own datasets. First, we are neither presenting an eco-driving proof of concept nor attempting to answer how to do eco-driving. Our goal is to analyze the impact of eco-driving to answer the question of whether to eco-drive. Such questions can not be answered by analyzing merely a few intersection scenarios. Unless we have a representative set of intersection scenarios, merely relying on a select intersection could lead to evaluation overfitting where the insights extracted from a few scenarios do not generalize to the class of scenarios (24; 57).
Furthermore, we must avoid overfitting to a specific region in our impact assessment (17). While there are popular intersection datasets available for Manhattan Island in New York (56), using them could lead to regional overfitting. Manhattan Island exhibits distinct spatial and temporal traffic dynamics due to its block-oriented road network. Insights drawn from an impact assessment solely based on these intersections would not be effective in generalizing to other cities or regions.
Another option is to use synthetic datasets, which are easier to generate (32). However, impact assessments based on these datasets are not particularly meaningful as they may not reflect the real-world intersection distributions and may contain oversimplified scenarios (16).
Thus, the fundamental requirement for modeling intersection scenarios lies in their diversity, encompassing variations at both the intersection and regional levels and making sure they are drawn from the real-world distribution. However, it is crucial to exercise caution when creating datasets, ensuring that the fused data originate from the same distribution. For instance, the IntD dataset (6) containing a set of intersections in Germany is popular within the research community, but combining it with U.S.-specific data like vehicle inflows may result in a distribution mismatch.
@@ 3.3 Methods for eco-driving control
Earlier studies on eco-driving have predominantly explored five control strategies for controlling vehicles. They include 1) dynamic programming, 2) heuristics, 3) trajectory optimization, 4) model-based reinforcement learning (model-based RL), and 5) model-free reinforcement learning (model-free RL). Each of these approaches exhibits distinct advantages and drawbacks. In Table 1, we offer a comparative analysis of these methods, focusing on five prevalent shortcomings.
@@
Dynamic programming[1,2] x
Heuristics[3] x x x
Trajectory optimization[4,5] x x *
Model-based RL[6,7] * x x
Model-free RL[8-10] x x x x
Table 1: A comparison of the shortcomings of five known methodologies for eco-driving control policy synthesis. The ’x’ indicates that the method does not demonstrate the shortcoming, a ’✓’ indicates that the method demonstrates the shortcoming, and a ’*’ indicates that the shortcoming may or may not present. [1](34), [2](10), [3](27), [4](26), [5](53), [6](30), [7](67), [8](25), [9](55), [10](3)
  • Optimality: Refers to the method’s ability to produce optimal results.
@@
  • Computational load: Indicates the inference time computational burden needed for obtaining solutions. This is crucial, considering that vehicles may not have powerful onboard computers. We note that low inference time computational burden does not necessarily mean low training time computational burden, as in the case of deep reinforcement learning. Given that training happens only once and does not involve the vehicle onboard computers, and inference happens continuously in the vehicle onboard computers, here we only focus on inference time computational burden. However, we note that our work focuses on not coming up with the most efficient inference time algorithm but rather assessing the impact of eco-driving.
  • Scale: Refers to the optimization problem’s scale, e.g., one intersection or two. Here, scale does not consider deployment scale directly, although some synergies may exist.
@@ 3.4 Reinforcement learning for vehicle-based traffic control
Contrary to conventional traffic control mechanisms such as traffic signals and stop signs, vehicle-based traffic control is increasingly being explored in recent research. Such vehicle-based control is referred to as Lagrangian traffic control and involves controlling a fleet of vehicles to achieve certain fleet-level (including both controlled and uncontrolled vehicles) objectives such as congestion management (51), reducing emissions (25), and fostering smoother traffic flow (58; 62). While various studies have presented proofs-of-concept, they still fall short of accurately modeling real-world scenarios (62; 59; 9) or go beyond simplified traffic problems (62).
Existing research often relies on simplified traffic models, such as simplified traffic blocks (62), to assess the benefits of Lagrangian traffic control methods. However, none of these studies explore the potential of large-scale implementations that could offer meaningful insights for practical decision-making. While simplified tasks are valuable for gaining preliminary insights into a method, they might not resonate strongly with policymakers and stakeholders due to the lack of rigorous analysis of the benefits at a scale that holds significance (14).